In the scattering of x-rays from a crystal of NaCl, it is assumed that the scattered x-rays undergo no change in wavelength. Show that this is a reasonable assumption by calculating the order of magnitude of the Compton wavelength for a sodium atom and a chlorine atom and comparing it with a typical x-ray wavelength of 0.10 nm.


1
Expert's answer
2021-12-21T12:16:35-0500

For the wavelength after scattering, we have:


"\lambda'=\lambda+\frac h{m_ec}(1-\cos\theta),"

for θ=0° "\lambda'=\lambda."